OpenStudy (anonymous):

Solve 5x2 – 11x + 2 = 0 using the Quadratic Formula. Quadratic Formula is ax^3+bx^2=cx right?

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

a=5 b=-11 c=2

OpenStudy (anonymous):

\[\frac{ 11\pm \sqrt{11^{2}-4(5)(2)} }{ 2(5) }\]

OpenStudy (anonymous):

\[\frac{ 11\pm \sqrt{121-40} }{ 10}\]

OpenStudy (anonymous):

\[\frac{ 11\pm \sqrt{81} }{ 10 }\]

OpenStudy (anonymous):

\[\frac{ 11\pm9 }{ 10 }\]

OpenStudy (anonymous):

\[\frac{ 2 }{ 10 }\]

OpenStudy (anonymous):

can you simplify that?

OpenStudy (anonymous):

1/5th sorry i was doing somethin

OpenStudy (anonymous):

@pgpilot326 did i do that right?

OpenStudy (anonymous):

messed up while plugging in the 11 should by negative before squared but itll come out positive

OpenStudy (anonymous):

plus or minus... you should get 2 answers.

OpenStudy (anonymous):

oh i thought that meant negative >.< my bad so one answer is \[\frac{ 1 }{ 5 }\]

OpenStudy (anonymous):

\[\frac{ 11+9 }{ 2 }\,\text {and }\frac{ 11-9 }{ 2 }\]

OpenStudy (anonymous):

Wait now im confused x.x

OpenStudy (anonymous):

\[\frac{ 20 }{ 2 }\]

OpenStudy (anonymous):

it also factors so you can always use that as a check!

OpenStudy (anonymous):

alright i messed up thought the +- meant negative so one is 20/2 can that be simplified?

OpenStudy (anonymous):

10/1

OpenStudy (anonymous):

sorry,\[\frac{ 11+9 }{ 10 }\, \text{ and } \frac{ 11-9 }{ 10 }\]

OpenStudy (anonymous):

so 2 and 1/5

OpenStudy (anonymous):

thanks @pgpilot326 >.<

OpenStudy (anonymous):

Oh i was right c':

OpenStudy (anonymous):

you're welcome!

OpenStudy (anonymous):

just wasnt 100% positive.